A pure sample of 6.50 × 106 mol of a pure molecular substance has a mass of 3.64 × 105 kg. Calculate the relative molecular mass of the substance and the number of molecules that would be contained in a pure sample of the substance of mass 7.50 kg.
Answer #1

Relative molecular mass can be calculated by dividing the mass of substance by the number of moles of substance

so

M = \frac{3.64*10^5}{6.50*10^6}

M = 5.6*10^{-2}Kg/mole

and to calculate the number of moles in 7.50 Kg we have to use the value of relative molecular mass calculated abow

n = \frac{7.50}{5.6*10^{-2}}moles

n = 1.3392857*10^2 moles

So number of moles in 7.50 Kg is 133.92857 moles